OpenStudy (anonymous):

The value of an uncirculated 'Mint State-65' 1950 Jefferson nickelminted in Denver is 3/2 the value 4 1945nickel minted in Philly in similar condition. Together the total value of the two coins is $20. What is the value of each coin. value of 1945 nickel $? value of 1950 nickel $?

OpenStudy (campbell_st):

is the denver coin \[\frac{3}{2} \times 4 Philly \]

OpenStudy (anonymous):

3/2 the value

OpenStudy (campbell_st):

ok... let P = the value of a Philly coin then the Denver coin is 3/2 x 4P = 6P so the value of a Philly and Denver coin is $20 the equation is P + 6P = 20 7P = 20 P = 20/7 P = 2.85714 so the Philly coin has a value of $2.86 the Denver coin is 6 x $2.85714 = 17.1428 so a Denver coin has a value of $17.14 and a Philly coin $2.86